### v1

# DefVINS: Visual–Inertial Odometry for Deformable Scenes

Abstract
Deformable scenes violate the rigidity assumptions underpinning classical visual–inertial odometry (VIO), often leading to over–fitting to local non–rigid motion or severe drift when deformation dominates visual parallax. We introduce DefVINS, a visual–inertial odometry framework that explicitly separates a rigid, IMU–anchored state from a non–rigid warp represented by an embedded deformation graph. The system is initialized using a standard VIO procedure that fixes gravity, velocity, and IMU biases, after which non–rigid degrees of freedom are activated progressively as the estimation becomes well conditioned. An observability analysis is included to characterize how inertial measurements constrain the rigid motion and render otherwise unobservable modes identifiable in the presence of deformation. This analysis motivates the use of IMU anchoring and informs a conditioning–based activation strategy that prevents ill–posed updates under poor excitation. Ablation studies demonstrate the benefits of combining inertial constraints with observability–aware deformation activation, resulting in improved robustness under non–rigid environments. Source will be released upon acceptance.
I Introduction
Simultaneous Localization and Mapping (SLAM), and its core state estimation component, Visual–Inertial Odometry (VIO), are foundational technologies driving modern mobile robotics and Augmented Reality (AR). Visual odometry and SLAM is a very active field of research with an abundance of applications in these fields, demonstrating remarkable maturity through various sensor configurations, including monocular [1] and stereo [2].
Among these, the fusion of vision with an Inertial Measurement Unit (IMU) stands out as a particularly robust solution. Adding an IMU helps dealing with untextured environments and rapid motions and makes roll and pitch directly observable [3]. The IMU, which provides high–frequency measurements of linear acceleration and angular velocity, effectively anchors the system’s short–term pose dynamics and provides crucial observability to the orientation axes unobservable in pure monocular vision. On the other hand, the camera complements the IMU with external referencing to the environment in 6 Degrees of Freedom (DoF), correcting the inevitable drift resulting from IMU sensor biases over time [4, 5]. State–of–the–art VIO systems, such as VINS–Mono [5] and OKVIS [6], have proven highly effective in static, rigid environments, establishing them as the de facto standard for ego–motion estimation.
However, the efficacy of classical VIO hinges entirely upon the fundamental assumption of scene rigidity [7]. This assumption is severely violated in scenarios featuring deformable objects such as human bodies or clothing. When the geometric model fails to account for non–rigid motion, VIO systems experience significant issues: the estimation often suffers from early over–fitting of the rigid model to local non–rigid motion, or worse, considerable drift when deformation parallax dominates inter–frame motion [8]. Consequently, achieving robust and accurate localization in truly dynamic and deformable scenes remains an open and critical research problem [9].
Beyond modeling challenges, the transition from rigid to deformable environments fundamentally alters the observability properties of the estimation problem. While classical analyses of VIO observability have shown that inertial sensing renders scale, gravity direction, and roll–pitch observable in rigid scenes, the impact of these properties in the presence of non–rigid motion remains largely underexplored. Deformation introduces additional latent degrees of freedom that can strongly couple with camera motion, leading to severe ill–conditioning and ambiguity when relying on visual information alone. In this context, inertial measurements play a critical role by anchoring the rigid body dynamics over short time horizons, effectively constraining the solution space and mitigating spurious correlations between ego–motion and non–rigid deformation. This work therefore includes an explicit observability analysis of the visual–inertial deformable odometry problem, highlighting how inertial terms significantly improve conditioning and provide essential structural constraints that are otherwise absent in purely visual formulations.
To address the instability and ill-conditioning inherent in extending VIO to non-rigid scenes, we introduce DefVINS: Robust Visual–Inertial Odometry for Deformable Scenes. Our framework utilizes an embedded deformation graph to explicitly model the non-rigid warp, carefully separating the rigid, IMU-anchored state from the scene’s non-rigid degrees of freedom.
II Related Work
Visual–Inertial Odometry (VIO) is a well–established paradigm for robust ego–motion estimation in rigid environments through the tight fusion of visual and inertial measurements [10, 11, 12]. Inertial sensing provides high–rate motion constraints that resolve metric scale, stabilize roll and pitch, and mitigate drift, while visual observations offer global referencing. The observability and consistency properties of such systems have been studied extensively under rigid–scene assumptions, showing that scale, gravity, velocity, and IMU biases become observable only under sufficient excitation and appropriate modeling choices [13, 14, 15]. These analyses, however, fundamentally assume that all observed features belong to a single rigid body.
When this assumption is violated, visual residuals become inconsistent and may bias the estimation or lead to divergence. Several approaches address dynamic scenes by detecting and suppressing independently moving elements through semantic segmentation or motion consistency checks [16]. While effective under moderate dynamics, these methods do not explicitly model deformation and instead discard measurements, limiting their applicability in scenes dominated by non–rigid motion. Other dynamic SLAM pipeline, such as Detect–SLAM, focus on handling multiple moving objects but remain purely visual and do not address continuous deformation or metric consistency, as summarized in Table I.
A complementary line of work explicitly models non–rigid scene geometry. DynamicFusion [17], DefSLAM [18], and NR–SLAM [19] represent deformation using embedded deformation graphs or learned warp fields, enabling compelling reconstructions. However, these systems rely exclusively on visual cues, lack inertial anchoring, and do not preserve metric scale, resulting in a strong coupling between camera motion and scene deformation and, consequently, ill–conditioned pose estimation under strong non–rigid motion.
Method Def. IMU Scale Obs. Analysis
ROVIO [20]
SVO2 [21]
VINS–Mono [11]
OKVIS [10]
ORB–SLAM3 [12]
DynaSLAM [16] (✗)
Detect–SLAM [22] (✗)
CUDA–SIFT–SLAM [23]
DynamicFusion [17]
DefSLAM [18]
NR–SLAM [19]
DefVINS (Ours)
TABLE I: Comparison with representative SLAM and VIO approaches. The table highlights whether each method explicitly models non–rigid deformation (Def.), uses inertial sensing for metric state estimation (IMU / Scale), and provides an explicit observability or conditioning analysis of the estimation problem (Obs. Analysis).
From a state–estimation perspective, prior work has extensively analyzed degeneracy, excitation, and observability in rigid VIO systems  [13, 15, 24, 25]. In deformable environments, however, a significant portion of the observed parallax may originate from scene motion rather than camera motion, effectively reducing rigid excitation and degrading conditioning in ways not captured by existing rigid–scene analyses. To date, no prior work explicitly characterizes how inertial measurements affect the observability and conditioning of visual–inertial estimation in the presence of non–rigid deformation.
Table I summarizes representative SLAM and visual–inertial approaches. To the best of our knowledge, no existing method jointly (i) integrates an explicit deformation model within a visual–inertial odometry pipeline, (ii) maintains a rigid, IMU–anchored reference to ensure metric consistency, and (iii) provides an explicit observability or conditioning analysis of the resulting non–rigid estimation problem. We emphasize that the Obs. Analysis column refers to an explicit characterization of observability or conditioning of the proposed estimation model, rather than the implicit observability properties of a given sensor configuration.
DefVINS addresses these limitations by embedding a deformation graph within a visual–inertial optimization framework and by explicitly analyzing the observability properties of the resulting system, demonstrating that inertial constraints significantly improve conditioning even under strong non–rigid motion.
III Notation and Preliminaries
Vectors are denoted by bold lowercase letters ($\mathbf{x}$), unit vectors by a check accent ($\check{\mathbf{x}}$), and matrices by bold uppercase letters ($\mathbf{A}$). Scalars are represented by lowercase letters ($a$). Rotation matrices $\mathtt{R}_{ab}\in\mathrm{SO}(3)$ denote the orientation of frame $b$ with respect to frame $a$; when the reference frame is the world frame $w$, the first subscript is omitted, i.e., $\mathtt{R}_{a}\doteq\mathtt{R}_{wa}$.
IMU kinematic propagation between time instants $t_{i}$ and $t_{j}$ follows the standard on-manifold preintegration formulation [26] and is given by
$\mathtt{R}_{j}$ $=\mathtt{R}_{i}\prod_{k=i}^{j-1}\mathrm{Exp}\!\left(\big(\tilde{\boldsymbol{\omega}}_{k}-\mathbf{b}^{g}_{k}-\boldsymbol{\eta}^{gd}_{k}\big)\,\Delta t\right),$ (1)
$\mathbf{v}_{j}$ $=\mathbf{v}_{i}+\mathbf{g}\,\Delta t_{ij}+\sum_{k=i}^{j-1}\mathtt{R}_{k}\big(\tilde{\mathbf{a}}_{k}-\mathbf{b}^{a}_{k}-\boldsymbol{\eta}^{ad}_{k}\big)\,\Delta t,$ (2)
$\mathbf{t}_{j}$ $=\mathbf{t}_{i}+\sum_{k=i}^{j-1}\Big[\mathbf{v}_{k}\Delta t+\tfrac{1}{2}\mathbf{g}\,\Delta t^{2}+\tfrac{1}{2}\mathtt{R}_{k}\big(\tilde{\mathbf{a}}_{k}-\mathbf{b}^{a}_{k}-\boldsymbol{\eta}^{ad}_{k}\big)\,\Delta t^{2}\Big],$ (3)
where $\mathbf{v}_{i}$ and $\mathbf{v}_{j}$ denote the linear velocities at times $t_{i}$ and $t_{j}$, $\mathbf{t}_{i}$ and $\mathbf{t}_{j}$ the corresponding positions, $\mathbf{g}$ the gravity vector, and $\Delta t$ the IMU sampling period, with $\Delta t_{ij}$ denoting the total integration interval. The measurements $\tilde{\boldsymbol{\omega}}_{k}$ and $\tilde{\mathbf{a}}_{k}$ correspond to the gyroscope and accelerometer readings at time $k$, affected by biases $\mathbf{b}^{g}_{k}$, $\mathbf{b}^{a}_{k}$ and additive noise terms $\boldsymbol{\eta}^{gd}_{k}$, $\boldsymbol{\eta}^{ad}_{k}$. During the initialization phase, both gyroscope and accelerometer biases are assumed to be approximately constant, i.e., $\mathbf{b}^{g}_{k}\approx\mathbf{b}^{g}$ and $\mathbf{b}^{a}_{k}\approx\mathbf{b}^{a}$.
IV Visual–Inertial Odometry
This section details the components of the proposed cost function. We first introduce the system state definition. We then describe the inertial and gravity-related terms that anchor the estimation to the IMU. Next, the visual reprojection factors derived from image measurements are presented. Finally, the non-rigid regularization terms, including elastic, viscous, and photometric components defined on the deformation graph, are introduced.
IV-A State definition
The system state is defined over a local temporal window spanning two consecutive keyframes, $\{t{-}1,t\}$. It comprises the rigid-body variables at both instants, namely orientation $\mathtt{R}_{t}\in\mathrm{SO}(3)$, velocity $\mathbf{v}_{t}\in\mathbb{R}^{3}$, and position $\mathbf{p}_{t}\in\mathbb{R}^{3}$, together with global inertial parameters, including the gyroscope and accelerometer biases $\mathbf{b}^{g},\mathbf{b}^{a}$ and the gravity direction $\hat{\mathbf{g}}\in\mathbb{S}^{2}$. We use the gravity direction representation introduced in [27], Sec. III–C. To account for non-rigid scene dynamics, the state is further augmented with a non-rigid substate $\boldsymbol{\xi}_{\mathrm{NR}}$, which collects the positions of all deformation nodes active within the window, stacked at times t1t{-}1 and $t$. Formally, the system state is given by
$\boldsymbol{\xi}=\big[\mathtt{R}_{t-1},\mathbf{v}_{t-1},\mathbf{p}_{t-1},\;\mathtt{R}_{t},\mathbf{v}_{t},\mathbf{p}_{t},\;\mathbf{b}^{g},\mathbf{b}^{a},\hat{\mathbf{g}},\;\boldsymbol{\xi}_{\mathrm{NR}}\big].$ (4)
In the following subsections, the corresponding residuals are developed.
IV-B IMU preintegration and gravity residuals
Inertial information is incorporated through IMU preintegration factors following the on-manifold formulation of [26]. Between two time instants $t_{i}$ and $t_{j}$, the relative motion is constrained by rotation, velocity, and position residuals,
$\mathbf{r}_{\Delta\mathtt{R}}$ $=\mathrm{Log}\!\left(\Delta\tilde{\mathtt{R}}_{t-1,t}^{\top}\mathtt{R}_{t-1}^{\top}\mathtt{R}_{t}\right),$ (5)
$\mathbf{r}_{\Delta\mathbf{v}}$ $=\mathtt{R}_{i}^{\top}\!\left(\mathbf{v}_{t}-\mathbf{v}_{t-1}-\mathbf{g}\Delta t_{t-1,t}\right)-\Delta\tilde{\mathbf{v}}_{t-1,t},$ (6)
$\mathbf{r}_{\Delta\mathbf{p}}=\mathtt{R}_{i}^{\top}\!\left(\mathbf{p}_{t}-\mathbf{p}_{t-1}-\mathbf{v}_{t-1}\Delta t_{t-1,t}-\tfrac{1}{2}\mathbf{g}\Delta t_{t-1,t}^{2}\right)-\Delta\tilde{\mathbf{p}}_{t-1,t}$ (7)
Each residual is weighted by its corresponding covariance matrix, obtained during the preintegration process as described in [27], Sec. III–A. Additionally, a gravity residual is introduced to enforce consistency between the estimated accelerations and the gravity direction,
$\mathbf{r}_{\mathbf{g}}=\left(\frac{\mathbf{v}_{t}-\mathbf{v}_{t-1}}{\Delta t_{t-1,t}}-\frac{\mathtt{R}_{t-1}\Delta\mathbf{v}_{t-1,t}}{\Delta t_{t-1,t}}\right)-\|\mathbf{g}\|\,\hat{\mathbf{g}}$, (8)
where $\hat{\mathbf{g}}\in\mathbb{S}^{2}$ and $\|\mathbf{g}\|=9.81~\mathrm{m/s}^{2}$. The contribution of the aforementioned residuals to the overall cost function is then defined as
$\mathcal{L}_{\mathrm{imu}}^{t}=\|\mathbf{r}_{\Delta\mathtt{R}}\|_{{\boldsymbol{\Sigma}}_{\Delta\boldsymbol{\phi}}}+\|\mathbf{r}_{\Delta\mathbf{v}}\|_{{\boldsymbol{\Sigma}}_{\Delta\mathbf{v}}}+\|\mathbf{r}_{\Delta\mathbf{p}}\|_{{\boldsymbol{\Sigma}}_{\Delta\mathbf{p}}}+\|\mathbf{r}_{\mathbf{g}}\|_{{\boldsymbol{\Sigma}}_{\mathbf{g}}}$ (9)
IV-C Visual reprojection term
The visual contribution is modeled through a standard geometric reprojection error, which enforces multi-view consistency between observed image features and their predicted projections under the estimated camera motion. Specifically, the visual cost at time $t$ is defined as
$\mathcal{L}_{\mathrm{vision}}^{t}=\sum_{k}\left\|\mathbf{z}_{k}-\pi\!\left(\mathtt{T}_{t},\mathbf{X}_{k}\right)\right\|_{\boldsymbol{\Sigma}_{k}},$ (10)
where $\mathbf{z}_{k}$ denotes the image measurement of feature $k$, $\mathbf{X}_{k}$ its corresponding 3D point, $\pi(\cdot)$ the camera projection model, and $\boldsymbol{\Sigma}_{k}$ the associated measurement covariance [28].
IV-D Non-rigid terms
Real environments may exhibit mild non-rigid motion, such as cloth or cable deformations. To model these effects without resorting to dense representations, a lightweight deformation graph built from long-term feature tracks is adopted. Each track $i$ defines a graph node with 3D position $\mathbf{x}_{i}^{t}$ at keyframe $t$. Two nodes are connected by an edge $(i,j)$ if their distance in the reference keyframe is below a spatial threshold. This provides a simple neighborhood structure that describes how nearby points of the scene relate to each other. The reference distance between nodes is defined as
$d_{ij}^{0}=\bigl\lVert\mathbf{x}_{i}^{0}-\mathbf{x}_{j}^{0}\bigr\rVert.$ (11)
The full non–rigid regularization is built up by means of specific terms which will be explained below.
$\mathcal{O}=\left[\begin{array}[]{cccccccccc}-\mathbf{J}_{r}^{-1}(\boldsymbol{\phi})\mathtt{R}_{j}^{\top}\mathtt{R}_{i}&\mathbf{0}&\mathbf{0}&\mathbf{J}_{r}^{-1}(\boldsymbol{\phi})&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\[3.01389pt] \left(\mathtt{R}_{i}^{\top}(\mathbf{v}_{j}-\mathbf{v}_{i}-\|\mathbf{g}\|\hat{\mathbf{g}}\Delta t_{ij})\right)^{\wedge}&-\mathtt{R}_{i}^{\top}&\mathbf{0}&\mathbf{0}&\mathtt{R}_{i}^{\top}&\mathbf{0}&\mathbf{0}&-\mathtt{R}_{i}^{\top}\Delta t_{ij}&-\|\mathbf{g}\|\mathtt{R}_{i}^{\top}\Delta t_{ij}&\mathbf{0}\\[3.01389pt] \left(\mathtt{R}_{i}^{\top}(\mathbf{p}_{j}-\mathbf{p}_{i}-\mathbf{v}_{i}\Delta t_{ij}-\tfrac{1}{2}\|\mathbf{g}\|\hat{\mathbf{g}}\Delta t_{ij}^{2})\right)^{\wedge}&-\mathtt{R}_{i}^{\top}\Delta t_{ij}&-\mathtt{R}_{i}^{\top}&\mathbf{0}&\mathbf{0}&\mathtt{R}_{i}^{\top}&\mathbf{0}&-\mathtt{R}_{i}^{\top}\tfrac{1}{2}\Delta t_{ij}^{2}&-\tfrac{1}{2}\|\mathbf{g}\|\mathtt{R}_{i}^{\top}\Delta t_{ij}^{2}&\mathbf{0}\\[3.01389pt] \mathbf{H}_{\mathtt{R}_{i}}^{\text{vis}}&\mathbf{0}&\mathbf{H}_{\mathbf{p}_{i}}^{\text{vis}}&\mathbf{H}_{\mathtt{R}_{j}}^{\text{vis}}&\mathbf{0}&\mathbf{H}_{\mathbf{p}_{j}}^{\text{vis}}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{H}_{\mathrm{NR}}^{\text{vis}}\\[3.01389pt] \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{H}_{\mathrm{NR}}^{\text{elas}}\\[3.01389pt] \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{H}_{\mathrm{NR}}^{\text{visc}}\\[3.01389pt] \mathbf{H}_{\mathtt{R}_{i}}^{\text{photo}}&\mathbf{0}&\mathbf{H}_{\mathbf{p}_{i}}^{\text{photo}}&\mathbf{H}_{\mathtt{R}_{j}}^{\text{photo}}&\mathbf{0}&\mathbf{H}_{\mathbf{p}_{j}}^{\text{photo}}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{H}_{\mathrm{NR}}^{\text{photo}}\\[3.01389pt] \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{I}_{3}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\[3.01389pt] \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{I}_{3}&\mathbf{0}&\mathbf{0}\\[3.01389pt] \mathbf{0}&-\tfrac{1}{\Delta t_{ij}}\mathbf{I}_{3}&\mathbf{0}&\mathbf{0}&\tfrac{1}{\Delta t_{ij}}\mathbf{I}_{3}&\mathbf{0}&\mathbf{0}&\mathbf{0}&-\|\mathbf{g}\|\,\mathbf{J}_{\mathbb{S}^{2}}&\mathbf{0}\par\end{array}\right]$
Figure 1: Structure of the full observability matrix $\mathcal{O}$. The matrix explicitly incorporates the accelerometer bias into the state. Each block row corresponds to the Jacobians of the different measurement and constraint terms, including IMU preintegration factors, visual residuals, non–rigid motion priors, and bias and gravity constraints. This structured formulation highlights the contribution of each sensing modality to the overall system observability.
Elastic term. The elastic prior prevents unrealistic stretching or compression of the graph. If the object bends slightly, nearby nodes may move, but their spacing should not deviate excessively from the reference configuration. This behaviour is encouraged by penalizing changes in pairwise distances:
$\mathcal{L}_{ij,\mathrm{elas}}^{t}=k\,\frac{\left(d_{ij}^{t}-d_{ij}^{0}\right)^{2}}{d_{ij}^{0}},\qquad d_{ij}^{t}=\bigl\lVert\mathbf{x}_{i}^{t}-\mathbf{x}_{j}^{t}\bigr\rVert.$ (12)
In intuitive terms, if nodes preserve their relative spacing the penalty is small; if they stretch or compress significantly, the penalty increases.
Viscous term. While the elastic term controls the shape, we also want to regularize the motion of neighboring nodes over time. Let $\mathbf{s}_{i}^{t}=\mathbf{x}_{i}^{t}-\mathbf{x}_{i}^{t-1}$ be the displacement of node $i$ between two consecutive keyframes. Following [29], we encourage nearby nodes to move in a similar way:
$\mathcal{L}_{ij,\mathrm{visc}}^{t}=b_{ij}\,\bigl\lVert\mathbf{s}_{i}^{t}-\mathbf{s}_{j}^{t}\bigr\rVert^{2},$ (13)
where proximity is encoded by the spatially decaying weights
$b_{ij}=\exp\!\left(-\frac{\left\lVert\mathbf{x}_{i}^{0}-\mathbf{x}_{j}^{0}\right\rVert^{2}}{2\sigma^{2}}\right).$ (14)
Nearby nodes therefore have a strong temporal coupling, promoting smooth and coherent deformations, while distant nodes influence each other only weakly.
Photometric term. Beyond geometric constraints, the formulation also exploits image intensity information. A semi-direct strategy is adopted, in which photometric data association is carried out on Shi–Tomasi features using the modified multi-scale Lucas–Kanade tracker introduced in [30]. When a deformation node is visible in both keyframes t1t{-}1 and $t$, its image projections are expected to correspond to pixels with similar intensities, following the classical brightness constancy assumption:
$\mathcal{L}_{i,\mathrm{photo}}^{t}=\Big(I^{t}(\mathbf{u}_{i}^{t})-\alpha_{i}I^{t-1}(\mathbf{u}_{i}^{t-1})+\beta_{i}\Big)^{2},$ (15)
where $\mathbf{u}_{i}^{t}$ is obtained by projecting $\mathbf{x}_{i}^{t}$ into the image and $I^{t}(\cdot)$ is evaluated via bilinear sampling. Also a local illumination invariance is achieved by computing local gain $\alpha_{i}$ and bias $\beta_{i}$ terms for each point. This term encourages nodes to move consistently with the apparent texture motion in the images.
Combining the previous components yields the full non-rigid regularization for keyframe pair (t1,t)(t{-}1,t):
$\mathcal{L}_{\mathrm{NR}}^{t}=\sum_{(i,j)\in\mathcal{E}}\!\Big(\mathcal{L}_{ij,\mathrm{elas}}^{t}+\mathcal{L}_{ij,\mathrm{visc}}^{t}\Big)+\sum_{i\in\mathcal{V}}\mathcal{L}_{i,\mathrm{photo}}^{t},$ (16)
with $\mathcal{V}$ the set of deformation nodes.
IV-E Overall cost function
The visual–inertial estimation problem is formulated as a nonlinear least-squares optimization over a fixed-size sliding window of keyframes. The window contains the most recent $N$ keyframes and defines the set of state variables over which all measurement constraints are applied. Within this window, the system jointly optimizes rigid-body motion, inertial parameters, and non-rigid deformation variables.
Let $\mathcal{W}=\{t_{0},t_{0}{+}1,\dots,t_{0}{+}N{-}1\}$ denote the active sliding window. As new keyframes are added, the oldest keyframe $t_{0}$ is removed from the window and its corresponding state variables are marginalized. This marginalization step condenses past information into a compact prior term $\mathcal{L}_{\mathrm{prior}}$, which preserves estimator consistency while keeping the computational complexity bounded.
Following [31], only consecutive temporal pairs $\{t{-}1,t\}$ with $t\in\mathcal{W}\setminus\{t_{0}\}$ are considered. For each pair, inertial preintegration, geometric reprojection, and non-rigid regularization terms are accumulated. The resulting objective function minimized by the system is given by
$\mathcal{L}=\sum_{t\in\mathcal{W}\setminus\{t_{0}\}}\Big(\mathcal{L}_{\mathrm{imu}}^{t}+\mathcal{L}_{\mathrm{vision}}^{t}+\lambda_{\mathrm{NR}}\,\mathcal{L}_{\mathrm{NR}}^{t}\Big)\,+\mathcal{L}_{prior},$ (17)
where $\mathcal{L}_{\mathrm{imu}}^{t}$, $\mathcal{L}_{\mathrm{vision}}^{t}$, and $\mathcal{L}_{\mathrm{NR}}^{t}$ are defined in (9), (10), and (16), respectively. The term $\mathcal{L}_{\mathrm{prior}}$ represents the prior induced by marginalization, while the scalar $\lambda_{\mathrm{NR}}$ balances the non-rigid regularization against the visual and inertial contributions.
IV-F Observability analysis
Following [27], observability is assessed from a discrete-time perspective by linearizing the measurement residuals over a finite temporal window. Although Lie derivatives are not explicitly computed, the resulting Jacobians capture the local sensitivity of the measurements to the state in a manner analogous to the continuous-time formulation. As shown in [32, 15], the rank of the stacked Jacobian matrix determines the locally observable directions and is critical for estimator consistency.
Within this framework, observability is analyzed over a two-keyframe window $\{t{-}1,t\}$, jointly considering inertial preintegration, geometric visual, and non-rigid deformation residuals acting on the state (4). The corresponding observability matrix $\mathcal{O}$ is obtained by stacking the Jacobians of all residuals with respect to the state vector $\boldsymbol{\xi}$, yielding
$\mathcal{O}=\begin{aligned} \Big[&\frac{\partial\mathbf{r}_{\Delta\mathtt{R}}^{t}}{\partial\boldsymbol{\xi}}^{\top},\frac{\partial\mathbf{r}_{\Delta\mathbf{v}}^{t}}{\partial\boldsymbol{\xi}}^{\top},\frac{\partial\mathbf{r}_{\Delta\mathbf{p}}^{t}}{\partial\boldsymbol{\xi}}^{\top},\frac{\partial\mathbf{r}_{\mathrm{vision}}^{t}}{\partial\boldsymbol{\xi}}^{\top},\frac{\partial\mathbf{r}_{\mathrm{elas}}^{t}}{\partial\boldsymbol{\xi}}^{\top},\\ &\hskip 16.38895pt\hskip 16.38895pt\frac{\partial\mathbf{r}_{\mathrm{visc}}^{t}}{\partial\boldsymbol{\xi}}^{\top},\frac{\partial\mathbf{r}_{\mathrm{photo}}^{t}}{\partial\boldsymbol{\xi}}^{\top},\frac{\partial\mathbf{r}_{\mathbf{b}^{g}}^{t}}{\partial\boldsymbol{\xi}}^{\top},\frac{\partial\mathbf{r}_{\mathbf{b}^{a}}^{t}}{\partial\boldsymbol{\xi}}^{\top},\frac{\partial\mathbf{r}_{\mathbf{g}}^{t}}{\partial\boldsymbol{\xi}}^{\top}\Big].\end{aligned}$ (18)
TABLE II: Results on the Drunkard’s Dataset. We report mean ATE RMSE, translational RPE, and average number of successfully tracked frames per scene. Each row corresponds to the average performance over all scenes at the same deformation level. Best result per metric in bold. Last row represents the mean for each column.
Seq. Deformation ATE RMSE [mm] RPE [mm] #Frames
ORB-SLAM3 NR-SLAM DefVINS ORB-SLAM3 NR-SLAM DefVINS ORB-SLAM3 NR-SLAM DefVINS
V-NR VI-R Full V-NR VI-R Full V-NR VI-R Full
L0 Low 6.0 5.4 9.2 7.1 6.8 1.1 1.0 2.2 1.9 1.2 1987 2061 2124 2169 2198
L1 Medium 19.4 11.6 17.1 13.2 9.4 2.1 2.0 3.1 2.4 2.0 1879 1968 2057 2096 2128
L2 Hard 42.3 19.5 27.4 21.1 14.3 3.2 3.0 4.1 3.4 3.1 1746 1842 1928 1986 2021
L3 Extreme 53.1 25.4 39.2 30.3 19.6 5.0 4.3 5.2 4.1 3.3 1612 1729 1814 1876 1919
Mean 30.2 15.5 23.2 17.9 12.5 2.85 2.6 3.7 3.0 2.4 1806 1900 1981 2032 2067
Based on the previously defined residuals, Fig. 1 illustrates the explicit block structure of the observability matrix $\mathcal{O}$. For clarity and readability, the analytical expressions of the nonzero Jacobian blocks are reported in the text rather than embedded in the figure. Geometric reprojection residuals contribute standard visual–inertial Jacobian blocks, namely $\mathbf{H}_{\mathtt{R}_{t}}^{\mathrm{vis}}=-\mathbf{J}_{\pi}\,\mathbf{p}_{\mathrm{cam}}^{\wedge}$, $\mathbf{H}_{\mathbf{p}_{t}}^{\mathrm{vis}}=-\mathbf{J}_{\pi}\mathtt{R}_{t}^{\top}$, and $\mathbf{H}_{\mathrm{NR}}^{\mathrm{vis}}=\mathbf{J}_{\pi}\mathtt{R}_{t}^{\top}$ for the deformation nodes. Non-rigid regularization terms introduce additional couplings among deformation variables: the elastic prior yields $\mathbf{H}_{\mathrm{NR}}^{\mathrm{elas}}\propto(\mathbf{x}_{i}^{t}-\mathbf{x}_{j}^{t})/d_{ij}^{t}$, the viscous term results in block structures of the form $\mathbf{I},-\mathbf{I},-\mathbf{I},\mathbf{I}]$, and the photometric term contributes $\mathbf{H}_{\mathrm{NR}}^{\mathrm{photo}}=\nabla I^{t}(\mathbf{u}_{i}^{t})\,\mathbf{J}_{\pi}\mathtt{R}_{t}^{\top}$, with analogous derivatives with respect to the pose variables. Together, these expressions define all nonzero blocks of $\mathcal{O}$.
A symbolic inspection of the resulting Jacobian structure shows that the formulation preserves the classical gauge freedoms of visual–inertial odometry while introducing additional degrees of freedom associated with the non-rigid deformation field. When the non-rigid variables $\boldsymbol{\xi}_{\mathrm{NR}}$ are ignored and metric depth is assumed, the rigid subsystem remains observable only up to a global $SE(3)$ transformation: absolute position and yaw are unobservable, whereas gravity magnitude and direction become observable under sufficiently rich accelerations and rotations [32]. In particular, the IMU velocity and position residuals $\mathbf{r}_{\Delta\mathbf{v}}^{t}$ and $\mathbf{r}_{\Delta\mathbf{p}}^{t}$ couple accelerometer bias and gravity, becoming rank-deficient under near-constant-velocity motion, while insufficient rotational excitation leaves yaw and gyroscope bias coupled in $\mathbf{r}_{\Delta\mathtt{R}}^{t}$.
The non-rigid substate $\boldsymbol{\xi}_{\mathrm{NR}}$ is constrained by the combined action of geometric reprojection, photometric consistency, and visco–elastic regularization. Visual and photometric terms anchor node motion to image evidence, whereas elastic and viscous priors regularize spatial and temporal variations, eliminating low-energy deformation modes that would otherwise be consistent with visual measurements alone. As a result, the rank of $\mathcal{O}$ increases in the non-rigid subspace.
Although inertial measurements do not directly observe the deformation field, they improve its observability indirectly by stabilizing the global rigid trajectory. Once gravity and inertial biases are sufficiently well conditioned, deformation nodes can no longer absorb errors in global pose or yaw, reducing their ability to mimic rigid drift. Consequently, sufficiently rich inertial excitation helps decouple rigid motion from non-rigid deformation and further constrains the remaining deformation modes.
In practice, the rank of $\mathcal{O}$ depends on both motion and deformation patterns. Trajectories with limited parallax or negligible rotation lead to weak observability of inertial quantities and non-rigid modes, whereas well-excited motions combined with visco–elastic regularization yield a well-conditioned system in which the rigid state is observable up to the expected gauge freedoms and the deformation field is significantly constrained.
V Experiments
All experiments were conducted on a desktop PC with an Intel Core i7-11700K (3.6 GHz, 64 GB RAM). The proposed method is implemented in C++ and optimized using Ceres [^1] http://ceres-solver.org/. Unless otherwise stated, all methods run in a single thread and wall-clock times are reported.
Performance is evaluated using standard trajectory-based metrics. We report Absolute Trajectory Error (ATE) and Absolute Rotation Error (ARE) to assess global accuracy, Relative Pose Error (RPE) in translation and rotation to measure drift, and the number of successfully tracked frames as an indicator of robustness. Comparisons are performed against ORB-SLAM3 [12] as a rigid SLAM baseline and NR-SLAM [19] as a non-rigid alternative.
We first evaluate the method on the Drunkard’s Dataset [33], which provides 19 synthetic sequences ($\times320$) with full 3D ground truth and four increasing deformation levels per scene, enabling controlled analysis. We then validate the approach on seven real RGB–D sequences ($8×\times4$80) recorded in an industrial setup following [34], using synchronized cameras, IMU, and ground-truth motion. The sequences (R0–R6) capture progressively stronger non-rigid deformations of a textured mandala cloth.
V-A Synthetic experiments
To evaluate the robustness of the proposed initialization strategies under controlled deformation conditions, synthetic experiments are conducted on the Drunkard’s Dataset. This dataset consists of synthetic RGB-D sequences exhibiting progressively increasing levels of non-rigid deformation, ranging from near-rigid motion to large-amplitude surface dynamics. Since inertial measurements are not provided, IMU data are synthetically generated by differentiating the ground-truth trajectories represented as B-splines, ensuring temporally smooth and physically consistent signals. Table II reports quantitative results on the Drunkard’s Dataset, which comprises 19 synthetic scenes evaluated at four increasing levels of deformation difficulty. Results are averaged across scenes at each deformation level to provide a compact yet representative comparison. In low-deformation scenarios (L0), all methods achieve very low errors, with differences in ATE remaining below approximately 35%. In this regime, the rigid visual–inertial baseline ORB-SLAM3 performs competitively with the proposed approaches, indicating that explicit non-rigid modeling is not strictly necessary when deformations are minimal. Nevertheless, the incorporation of inertial measurements already provides measurable benefits: compared to the visual-only DefVINS configuration (V-NR), the rigid visual–inertial variant (VI-R) reduces ATE by roughly 20%. As deformation increases to moderate levels (L1), performance differences become more pronounced. While ORB-SLAM3 exhibits a substantial increase in ATE relative to L0, both NR-SLAM and DefVINS maintain significantly lower errors. In particular, the full DefVINS formulation achieves an ATE reduction of approximately 30% with respect to ORB-SLAM3 and nearly 45% compared to the visual-only baseline.
Improvements in RPE remain more moderate (around 20–30%), suggesting that inertial constraints primarily contribute to stabilizing short-term motion estimates. In harder deformation regimes (L2), rigid motion assumptions further degrade estimation accuracy. Relative to ORB-SLAM3, the full DefVINS formulation reduces ATE by approximately 35–40% and RPE by about 25%. Moreover, when compared to the rigid DefVINS variant (VI-R), the inclusion of explicit non-rigid regularization yields an additional ATE reduction of roughly 30%, highlighting the importance of modeling deformation dynamics beyond inertial stabilization alone. In the extreme deformation case (L3), the limitations of rigid models become most evident. The full DefVINS formulation achieves an ATE reduction of approximately 40–45% with respect to ORB-SLAM3 and nearly 50% compared to the visual-only configuration, while also reducing RPE by about 40%. Although all methods experience a reduction in the number of successfully tracked frames as deformation severity increases, DefVINS consistently maintains longer trajectories, indicating improved robustness to severe non-rigid motion.
Figure 2: Illustrative observability analysis under synthetic conditions. Evolution of the conditioning score $\log_{10}(\rho_{k})$ as a function of the number of stacked keyframe pairs $k$. Inertial sensing and non-rigid regularization significantly improve numerical conditioning, yielding well-observed directions with only a few frames.
Overall, results on the Drunkard’s Dataset confirm that inertial sensing and explicit non-rigid regularization provide complementary benefits. While inertial constraints primarily improve local motion consistency, non-rigid modeling is essential to recover globally accurate trajectories as deformation amplitude increases, even under idealized synthetic conditions.
TABLE III: Comparison of visual–inertial odometry methods on our real deformable sequences. We report ATE RMSE, translational RPE, and number of successfully tracked frames. Best result per metric in bold. Last row represents the mean for each column.
Seq. Deformation ATE RMSE [mm] RPE [mm] #Frames
ORB-SLAM3 NR-SLAM DefVINS ORB-SLAM3 NR-SLAM DefVINS ORB-SLAM3 NR-SLAM DefVINS
V-NR VI-R Full V-NR VI-R Full V-NR VI-R Full
R0 Low 6.9 7.1 10.8 8.9 8.1 1.2 1.4 1.9 1.6 1.5 1810 1804 1805 1802 1804
R1 Low 7.5 8.6 13.2 10.6 9.0 1.4 1.6 2.1 1.8 1.6 1684 1743 1751 1743 1770
R2 Medium 15.3 10.5 19.6 15.1 10.2 2.3 2.2 2.7 2.5 2.0 1589 1658 1706 1742 1768
R3 Medium 27.6 17.9 26.4 19.9 10.8 3.5 3.0 3.1 2.5 2.1 1496 1592 1651 1688 1719
R4 Medium 48.1 19.4 30.8 24.9 11.4 4.6 3.4 3.1 2.0 1.9 1387 1504 1568 1603 1736
R5 High 71.4 39.8 44.5 33.2 15.6 6.1 4.5 4.0 2.9 2.4 1194 1542 1623 1668 1706
R6 High 95.8 57.2 60.8 41.0 19.8 7.8 5.2 4.7 3.5 3.0 982 1476 1558 1604 1641
Mean 39.0 22.9 29.4 22.0 12.1 3.84 3.04 3.09 2.40 2.07 1449 1617 1666 1693 1735
Figure 3: Qualitative comparison of DefVINS operating modes and baselines under R4 sequence. a) Rigid VI SLAM (ORB-SLAM3): a representative state-of-the-art rigid visual–inertial system, which may suffer from tracking loss and relocalization in deformable scenes. b) Non-rigid SLAM (NR-SLAM): a representative state-of-the-art for non-rigid environments, which also suffer from tracking loss and relocalization due to its focus on medical datasets. c) DefVINS-V (visual-only, non-rigid): absence of inertial constraints leads to accumulated drift, particularly during turning motions. b) DefVINS-VI (rigid): introducing inertial sensing stabilizes rotational estimates, but assuming scene rigidity limits performance under deformation. e) DefVINS (full): by jointly enabling inertial sensing and explicit non-rigid modeling, the proposed system achieves the highest global consistency and lowest trajectory error. Dashed line represent the ground-truth while colored line represent the camera trajectory.
V-B Observability analysis
To complement the quantitative evaluation, we present an illustrative analysis of numerical observability under controlled synthetic conditions. A set of short motion segments with rich 6-DoF excitation is selected from the synthetic dataset. Specifically, we extract five non-overlapping segments of ten consecutive keyframes each, characterized by non-collinear translations and rotations about multiple axes. For each segment, residual Jacobians are generated and stacked over $k$ consecutive keyframe pairs to form the observability matrix $\mathcal{O}$. The reported results correspond to the average conditioning score across the selected segments. Numerical observability is quantified using the conditioning score $\rho_{k}={\sigma_{\min}}/{\sigma_{\max}}$, where $\sigma_{\min}$ and $\sigma_{\max}$ denote the smallest and largest singular values of $\mathcal{O}$, respectively. Fig. 2 shows $\log_{10}(\rho_{k})$ as a function of the number of stacked keyframe pairs for the visual-only, rigid visual–inertial, and full formulations. The numerical conditioning evolves very differently for the aforementioned configurations. Inertial constraints lift near-null modes associated with gravity and biases, and non-rigid regularization prevents deformation variables from absorbing rigid-body drift. As a result, the full formulation becomes well conditioned with only a few frames, explaining the improved robustness observed in the synthetic experiments.
V-C Realistic experiments
The synthetic evaluation is complemented with a set of realistic experiments aimed at assessing the behavior of the proposed non-rigid regularization under natural image noise, real sensor motion, and uncontrolled surface deformations. To this end, a dedicated dataset composed of seven RGB-D sequences is employed, each captured on a textured deformable surface and exhibiting a different degree of non-rigid motion. The sequences are denoted as R0–R6, where R0 corresponds to near-rigid motion and R6 represents the most severe deformation. This gradual increase in deformation enables a systematic evaluation of the stability of the proposed visco–elastic prior under progressively more challenging conditions. Table III summarizes the quantitative performance of all evaluated methods on the real deformable sequences R0–R6. Several consistent trends can be observed across deformation regimes.
For the nearly rigid sequences (R0–R1), the rigid visual–inertial baseline ORB-SLAM3 achieves the lowest ATE and RPE values, outperforming the deformable models by approximately 20–30% in ATE. This confirms that classical rigid visual–inertial approaches remain highly effective when surface deformation is negligible. In this regime, all DefVINS variants are able to track almost the full sequence length, with only marginal accuracy degradation. As the degree of deformation increases to moderate levels (R2–R3), rigid assumptions progressively break down. Although ORB-SLAM3 and NR-SLAM still maintain tracking, both exhibit a marked increase in ATE. In contrast, incorporating inertial measurements within the DefVINS framework (VI-R) reduces ATE by approximately 25–30% with respect to the visual-only configuration (V-NR), while also lowering RPE by around 15–20%. This highlights the stabilizing effect of inertial constraints on rotational estimation and short-term motion consistency. NR-SLAM consistently yields intermediate performance, improving upon rigid ORB-SLAM3 while remaining less robust than deformable formulations. For medium-to-high deformation scenarios (R4–R6), the limitations of rigid motion models become evident. Relative to ORB-SLAM3, the full DefVINS formulation reduces ATE by approximately 75% on R4, around 80% on R5, and close to 80% on R6. At the same time, DefVINS (Full) consistently tracks a substantially larger fraction of frames (around 85–95%) compared to ORB-SLAM3, whose tracking coverage drops below 50% in R4 and below 20% in R6. While inertial sensing alone partially mitigates tracking degradation, only the joint combination of inertial constraints and explicit non-rigid regularization preserves both accuracy and robustness under strong deformation.
Overall, these results demonstrate that inertial sensing and non-rigid modeling play complementary roles. Inertial measurements primarily stabilize local motion and reduce short-term drift, whereas explicit non-rigid regularization is essential to maintain global consistency and long-term tracking in highly deformable environments.
To assess the contribution of each system component, the proposed framework is evaluated under multiple operating modes obtained by selectively enabling inertial sensing and non-rigid modeling. Fig. 3 qualitatively compares the resulting camera trajectories for a visual-only configuration (DefVINS-V), a rigid visual–inertial setup without non-rigid regularization (DefVINS-VI), and the full DefVINS model, which jointly exploits inertial constraints and non-rigid motion modeling. Results are also compared against ORB-SLAM3 and NR-SLAM. This ablation study highlights the complementary roles of inertial sensing and non-rigid regularization, and exposes the limitations of rigid motion assumptions in deformable scenes.
To illustrate the estimated surface behavior, Fig. 4 shows the deformation graph for a representative sequence. The graph reveals spatially varying deformation, with larger displacements in the lower-left region and comparatively stable behavior elsewhere, demonstrating the ability of the proposed visco–elastic regularization to capture localized deformations while maintaining global surface coherence.
Figure 4: Deformation graph on sequence R4. Green and red edges denote the graph at times t1t\!-\!1 and $t$, respectively. Their differences indicate medium-to-high non-rigid deformation, with stronger effects in the lower-left region.
VI Conclusions
This paper introduced DefVINS, an observability-gated visual–inertial odometry framework for deformable scenes that decouples a rigid, IMU-anchored state from progressively activated non-rigid deformation parameters. Extensive evaluations on both synthetic and real datasets demonstrate that the proposed formulation provides accurate and stable state estimation across a wide range of deformation regimes. By combining IMU anchoring with conditioning-aware activation of deformation degrees of freedom, DefVINS avoids early over-fitting and catastrophic drift, constituting a suitable and reliable solution for visual–inertial odometry in the presence of non-rigid scene dynamics.